OpenStudy (anonymous):

Find the required measurment. Find the length of an arc of 40° in a circle with an 8 inch radius. (Use 3.14 for and round to the nearest hundredths.)

OpenStudy (anonymous):

length=2(40/360)(3.14)(8) solve it to get the answer

OpenStudy (anonymous):

then round the answer upto two decimal place

OpenStudy (anonymous):

2*pi*r =2*3.14*8 =50.24 40 degree = 40/360 * 50.24 = 1/9*50.24 =50.24/9 =5.58

OpenStudy (anonymous):

mathematicalpro gj......

OpenStudy (anonymous):

oh thnx hashir